# One Model to Train them All: Hierarchical Self-Distillation for Enhanced Early Layer Embeddings

### v1

@@ Abstract
Deploying language models often requires handling model size vs. performance trade-offs to satisfy downstream latency constraints while preserving the model’s usefulness. Model distillation is commonly employed to reduce model size while maintaining acceptable performance. However, distillation can be inefficient since it involves multiple training steps. In this work, we introduce ModularStarEncoder, a modular multi-exit encoder with 1B parameters, useful for multiple tasks within the scope of code retrieval. ModularStarEncoder is trained with a novel self-distillation mechanism that significantly improves lower-layer representations—allowing different portions of the model to be used while still maintaining a good trade-off in terms of performance. Our architecture focuses on enhancing text-to-code and code-to-code search by systematically capturing syntactic and semantic structures across multiple levels of representation. Specific encoder layers are targeted as exit heads, allowing higher layers to guide earlier layers during training. This self-distillation effect improves intermediate representations, increasing retrieval recall at no extra training cost. In addition to the multi-exit scheme, our approach integrates a repository-level contextual loss that maximally utilizes the training context window, further enhancing the learned representations. We also release a new dataset constructed via code translation, seamlessly expanding traditional text-to-code benchmarks with code-to-code pairs across diverse programming languages. Experimental results highlight the benefits of self-distillation through multi-exit supervision.
One Model to Train them All: Hierarchical Self-Distillation for Enhanced
Early Layer Embeddings
Andrea Gurioli1, Federico Pennino1, Joao Monteiro2 Maurizio Gabbrielli1 1University of Bologna, 2Autodesk Correspondence: andrea.gurioli5@unibo.it, joao.monteiro@autodesk.com
@@ 1 Introduction
Large language models (LLMs) have significantly impacted the field of natural language processing, demonstrating remarkable performance across various applications (Niu et al., 2023). However, the amount of computation required to operate state-of-the-art models poses significant challenges for the large-scale deployment of these models.
To mitigate these challenges, the research community has explored several model strategies to reduce the operational cost of LLMs without sacrificing their effectiveness. A prominent technique in model compression is quantization (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024), which involves the reduction of numerical precision in the model’s parameters. Quantization effectively decreases memory requirements and enhances inference speed, facilitating the deployment of large language models in resource-constrained environments. Concurrently, knowledge distillation has emerged as a powerful technique whereby a smaller “student” model is trained to emulate the behavior of a larger “teacher” model, as evidenced by works such as DistilBERT (Sanh et al., 2019) and TinyBERT (Jiao et al., 2019). Additionally, pruning methods selectively eliminate less influential weights or neurons, further reducing model complexity and aiming to preserve performance (Han et al., 2015).
Recent efforts have increasingly focused on developing efficient architectures requiring fewer parameters. Model families such as LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a paradigm shift towards smaller, more accessible architectures. These model families are deployed at various resolutions—ranging from lightweight variants optimized for heavily resource-constrained environments to larger versions that retain competitive performance.
@@
Figure 1: Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36). Each exit head predicts an output embedding and adds a “layer loss,” contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$.
In parallel, advancements in dynamic inference strategies have introduced mechanisms that further optimize computational efficiency. Techniques like multi-exit networks enable early predictions at intermediate layers, reducing unnecessary computations. For instance, early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) dynamically balance computation and accuracy by allowing predictions before full model execution. Similarly, Matryoshka representation learning (Kusupati et al., 2022) extends this idea to embeddings, introducing a loss function that yields multi-granular representations. This approach allows downstream tasks to adjust computational complexity by pruning embedding dimensionality, further contributing to efficient model deployment.
Building on these principles, we propose ModularStarEncoder, a modular multi-exit encoder architecture that integrates a novel intra-model self-distillation mechanism. In our design, specific intermediate layers are supervised by both the primary task loss and auxiliary distillation losses on specific exit heads, encouraging lower layers to learn better representations by mimicking the outputs of higher layers. We apply a shared embedding head comprising a masked language modeling head and an in-context classification head across a chosen subset of layers. We then fine-tuned the model with different projection heads for each exit point. We reached state-of-the-art results on multiple retrieval tasks (such as code-to-code and text-to-code), fine-tuning one single modular model that can be sliced depending on the end-user computational constraints.
@@
  • We introduce a self-distillation framework that enables training multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability. We believe this approach can significantly affect LLM training pipelines that depend on multiple model distillations.
  • We train and release ModularStarEncoder, which consists of a pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned encoder: The former is a modular pre-trained encoder with up to 1 billion parameters and five exit points, allowing users to perform multiple exit fine-tuning depending on downstream tasks. The latter is a fine-tuned encoder for various retrieval tasks. We allow the user to choose either the entire model with 1 billion parameters or a model size that fits their memory and computational limitations.
@@
  • We release SynthCode2Code2NL a new dataset [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCode2Code2NL constructed via code translation, expanding popular text-to-code datasets across diverse programming languages with code-to-code pairs. SynthCode2Code2NL comprises 1 071 367 triplets of natural language-code-code.
2 Methodology
2.1 Dataset
Table 1: SynthCode2Code2NL details: Average character count and sample size per language for the CodeSearchNet dataset and the synthesized portion obtained through translation.
Language CSN samples CSN avg. char Synth. samples Synth. avg. char
English 1 071 367 180 - -
PHP 280 706 514 116 967 579
Python 274 454 474 117 374 518
Go 234 089 350 124 125 541
Java 282 118 505 116 098 707
C++ - - 141 956 938
Ruby - - 158 494 456
C - - 136 365 1029
JavaScript - - 159 988 557
Table 2: Hyperparameters for Architecture, Pre-training, and Fine-tuning
Architecture
Hyperparameter Value
Model size 1B parameters
Precision bfloat16
Hidden layers 36
Attention heads 16
Hidden dimensionality 1024
Positional encoding RoPE ($\theta=10^{-6}$)
Context length 2048
Attention mechanism Grouped-Query Attention
Attention pattern Bi-directional
Pre-training
Batch size 3.99M tokens
Pretraining steps 245 000
Pretraining Tokens 1T
Loss function MLM + In-Context loss
Multi-layer loss yes
Optimizer AdamW
Weight decay 1e-1
Initial learning rate 6.24e-4
Learning rate schedule Multi-step
Warmup steps 4000
Fine-tuning
Dataset size 635.404 samples
Fine-tuning steps 20 000
Loss function CLIP loss
Multi-layer loss yes
Batch size 2048
Learning rate 1.0e-5
Temperature parameter 10.0
Hardware (Pre-training + fine-tuning)
GPUs 512 NVIDIA Ampere (64GB)
Overall Training hours 450 000
In the pre-training phase, we leveraged The Stack V2 Lozhkov et al. (2024), a large open-source code dataset structured by repository.
For the fine-tuning stage, we created SynthCode2Code2NL, a dataset that supports text-to-code and code-to-code search. Using the popular CodeSearchNet Husain et al. (2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript. We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig. 2). During code translation, we choose the token with the highest probability as output (greedy search) to prevent semantic discrepancies.
This process yielded pairs of code snippets in distinct languages tied to the same natural language description. As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages. SynthCode2Code2NL contains 1 071 367 samples where, in the first code column, we directly processed code snippets from CodeSearchNet, including Python, Java, PHP, and Go. The third column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets. After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments. Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port). During the preprocessing phase of SynthCode2Code2NL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column. During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams. Table 1 shows the average number of characters per language in SynthCode2Code2NL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in Limitations. Appendix A provides examples of code translation.
@@ 2.1 Dataset
Figure 3: On the left side the illustration of the in-context loss framework, where samples from different repositories are concatenated. Positive examples share the same repository context, while negative examples come from different repositories. On the right side, in-context loss framework pseudocode.
2.2 Architecture
We updated the first version of StarEncoder (Li et al., 2023) by enabling longer code snippets (up to 2 048 tokens as context length), increasing the model size from $\approx$125M to $\approx$1B parameters and utilizing state-of-the-art methodologies (Warner et al., 2024; Lozhkov et al., 2024) resulting in ModularStarEncoder.
We built ModularStarEncoder on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model. We reduced its size from 15B to 1B parameters. Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads. ModularStarEncoder relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1024 with an intermediate size of 12 288.
We followed Devlin et al. (2019) and replaced the causal self-attention in StarCoder-2 with bidirectional self-attention. Aiming for modularity, we also replaced sliding window attention with full attention. This step was taken to avoid the receptive field phenomenon of sliding window mechanisms (Zhu et al., 2021). Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference. Table 2 summarizes the architectural details.
2.3 Pre-training
We pre-trained ModularStarEncoder with a batch size of 3.99M tokens for 245 000 training steps, processing $\approx$1T tokens. We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs using the Leonardo supercomputer (Turisini et al., 2023), requiring 450 000 GPU working hours.
To enable both token-level and snippet-level embeddings after pre-training, we employed a multi-objective pre-training strategy that combined two losses, as detailed in section 2.3.1 and section 2.3.2. The pre-training was performed on TheStackV2, whose context length analysis revealed an average of $\approx 630$ tokens per code snippet. As described in section 2.3.1, we concatenated multiple snippets to facilitate our multi-loss methodology, allowing our in-context classification loss to expand the average context window to $\approx 1300$ tokens, reaching the maximum context length 20% of the time.
We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to 1e-6, and a weight decay of 1e-1. We initialized the learning rate at 6.24e-4 and decreased it using a multi-step learning rate schduler Bi et al. (2024) with 4 000 warmup steps. The learning rate was reduced at 120 000, 185 000, 220 000, 230 000, and 240 000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001. Table 2 summarizes the hyperparameters for architecture, pre-training, and fine-tuning.
2.3.1 Masked Language Modeling and In-Context Classification
The training objectives of BERT (Feng et al., 2020), specifically Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), have become a de facto standard. However, The NSP loss constrains the context window length to the sentence length, leading to too many padding tokens and redundant computation (Zeng et al., 2022), and has been shown to not yield significant benefits after fine-tuning (Warner et al., 2024; Aroca-Ouellette and Rudzicz, 2020). Given that the average number of tokens per data sample in Stack v2 is 630, a large context window of 2048 results in substantial padding, making long-context training inefficient. While Wang et al. (2023) demonstrated the advantages of training LLMs with multiple objectives, we revisited the NSP loss and introduced an in-context classification (ICC) objective. We hypothesize that predicting whether multiple code snippets belong to the same context (in our case, the same repository) can enhance semantic search performance while allowing efficient concatenation of multiple code fragments. Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
In $\mathcal{L}_{MLM}$, a certain percentage of tokens are randomly masked and predicted using a classification head. Following Zhang et al. (2024), we adopt a 15% masking rate with the standard 80-10-10 token replacement strategy Devlin et al. (2019). The secondary objective, $\mathcal{L}_{ICC}$, determines whether randomly concatenated inputs (separated by a $<SEP>$ token) originate from the same repository (see fig. 3). Each concatenated sample has a 50% probability of containing source code from different repositories. This approach increases input density—reducing padding by expanding the average input length from 630 to 1 300 tokens—and potentially enhances cross-language understanding. Since repositories are inherently modular and often contain files written in multiple languages, learning from repository-level context may improve inter-language generalization.
@@ 2.3.1 Masked Language Modeling and In-Context Classification
Table 3: Performance of different models on text-to-code with CodeSearchNet using codeXGLUE. We reported the results presented in codet5plus, unixcoder and modernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024).
@@
2.3.2 Multi-layer Loss
@@
To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.3.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index. The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied. The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set. This approach allows for flexible model deployment, enabling adaptive layer pruning while maintaining performance trade-offs.
2.4 Fine-tuning
Following Su et al. (2023), we fine-tune a single model for both text-to-code and code-to-code retrieval using instruction prompting. The optimization objective combines CLIP loss (Radford et al., 2021) with a multi-layer loss (details in 2.3.2).

### v2

# MoSE: Hierarchical Self-Distillation Enhances Early Layer Embeddings

@@ Abstract
Deploying language models often requires navigating accuracy vs. performance trade-offs to meet latency constraints while preserving utility. Traditional model distillation reduces size but incurs substantial costs through training separate models. We introduce ModularStarEncoder (MoSE), a 1-billion-parameter multi-exit encoder for code retrieval and classification that employs a novel Self-Distillation mechanism. This approach significantly enhances lower-layer representations, enabling flexible deployment of different model portions with favorable performance trade-offs. Our architecture improves text-to-code and code-to-code search by targeting specific encoder layers as exit heads, where higher layers guide earlier ones during training—improving intermediate representations at minimal additional cost. We further enhance MoSE with a repository-level contextual loss that maximizes training context window utilization. Additionally, we release a new dataset created through code translation that extends text-to-code benchmarks with cross-language code-to-code pairs. Evaluations demonstrate the effectiveness of Self-Distillation as a principled approach to trading inference cost for accuracy across various code understanding tasks.
@@ 1 Introduction
Large language models (LLMs) have significantly impacted natural language processing (Niu et al., 2023), but their computational demands pose substantial deployment challenges. The field has responded with various efficiency strategies: quantization techniques reduce numerical precision (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024); knowledge distillation trains smaller “student” models to emulate larger “teacher” models (Sanh et al., 2019; Jiao et al., 2019); and pruning methods eliminate less influential weights (Han et al., 2015). Concurrently, model families like LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a shift toward more efficient architectures at varying parameter scales.
Dynamic inference approaches further optimize efficiency through mechanisms like multi-exit networks. Early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) balance computation and accuracy by allowing predictions at intermediate layers, while Matryoshka representation learning (Kusupati et al., 2022) enables adjustment of computational complexity through embedding dimensionality pruning. These approaches maintain performance while reducing computational demands in resource-constrained environments. Furthermore, insights from the SimCLR framework (Chen et al., 2020) revealed that the final layer of a model does not necessarily produce the most semantically meaningful representations. Valeriani et al. (2023) analyzed the geometric evolution of hidden representations within large transformer models across layers, outlining that intermediate layers are identified as holding the most semantically rich representations for downstream tasks.
[…]
@@
Figure 1: (a) Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36). Each exit head predicts an output embedding and adds a layer loss, contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$. (b) Computational cost (GFLOPs) vs. performances trade-off at different exit layers over CodeSearchNet dataset (text-to-code in avg. MRR) and POJ104 (code-to-code in mAP). Despite a reduction of approximately 90% floating point operations from layer 36 to layer 4, MRR performance only drops by 6.4% in absolute terms. For POJ104, our best results are observed in the initial layers.
Building on these principles, we introduce ModularStarEncoder (MoSE), a modular multi-exit encoder built on top of the StarCoder-2 (Lozhkov et al., 2024) architecture, with 1 billion parameters that incorporate intra-model self-distillation. As illustrated in Figure 1(a), our architecture applies training objectives at multiple exit points throughout the network (layers 4, 9, 18, 27, and 36), with each exit contributing a weighted loss component to the overall objective. This approach encourages lower layers to learn better representations by mimicking higher-layer outputs. Figure 1(b) demonstrates the resulting accuracy-computation trade-off of encoders trained under this framework: despite a 90% reduction in floating point operations from layer 36 to layer 4, performance over text-to-code retrieval drops by only 6.4% in absolute terms, while the model maintains the best performances for code-to-code retrieval in very early layers.
Notably, the exit point with the best performance isn’t necessarily the final layer, as noted in previous findings due to Chen et al. (2020); Valeriani et al. (2023). This offers both deployment flexibility and a natural model selection mechanism for specific tasks. Users can select which portion of the model to use during inference, starting from the first exit point at 160M parameters. By allowing users to select an earlier exit point, MoSE provides a direct mechanism to decrease computational demands. This flexibility is crucial as smaller models generally require less energy and memory for inference, leading to more sustainable deployment options (Samsi et al., 2023).
We fine-tuned MoSE on SynthCoNL (cf. Sec. 5 for details) for text-to-code and code-to-code retrieval, and on BigCloneBench (Svajlenko & Roy, 2015) for code clone detection, achieving state-of-the-art results among open models while maintaining deployment modularity.
@@
    A self-distillation methodology that trains multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability—an approach that could significantly impact LLM training pipelines dependent on multiple model distillations.
  2. 2.
    MoSE, pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned, with up to 1 billion parameters and five exit points, allowing users to select model size based on memory and computational constraints.
  3. 3.
@@
    SynthCoNL [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCoNL, a new dataset of 1,071,367 natural language-code-code triplets constructed via code translation, expanding text-to-code benchmarks with cross-language code-to-code pairs.
@@
Figure 2: The illustration on the left depicts the in-context loss framework, where samples from various repositories are concatenated. Positive examples originate from the same repository context, whereas negative examples are sourced from different repositories. To enable the model’s use of FlashAttention V2, we applied left padding and positioned the CLS token at the end of the sentence. On the right side, you’ll find the pseudocode for the in-context loss framework.
2 Self-Distillation and In-Context Classification
As different layers of transformer models capture varying degrees of semantic richness, often with intermediate layers proving most effective for specific tasks (Bolya et al., 2025), we design our pre-training strategy to support modularity and ease downstream task adaptability. Rather than treating the deepest layer as the sole source of meaningful representations, we aim to enhance representations throughout the network.
We hypothesize that applying supervision at multiple depths, not just at the output—encourages more robust and semantically rich intermediate representations. To this end, we combine multiple training objectives, each contributing to more versatile and generalizable embeddings (Wang et al., 2023). While we retain the widely adopted Masked Language Modeling (MLM) loss Feng et al. (2020), we discard the traditional Next Sentence Prediction (NSP) objective. NSP has been shown to offer minimal benefit after fine-tuning (Liu et al., 2019), and it imposes a rigid sentence-level structure that leads to inefficient use of context windows, especially for long-form code inputs. In its place, we introduce an In-Context Classification (ICC) loss (see section 2.1) that increases input density by reducing padding and fosters better chunk-level comprehension, essential for tasks like semantic retrieval or classification.
By integrating these two complementary objectives, MLM and ICC, and distributing their supervision across multiple layers, forming the basis of our Self-Distillation mechanism (see section 2.2), our training approach improves internal representation capabilities and flexibility, ultimately yielding a model better suited for modular deployment and early exiting.
2.1 Masked Language Modeling and In-Context Classification
We revisited the NSP loss and introduced an In-Context Classification (ICC) objective. We hypothesize that predicting whether multiple code snippets belong to the same context (in our case, the same repository) can enhance semantic search performance while allowing efficient concatenation of multiple code fragments. Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
[…]
@@ 2.1 Masked Language Modeling and In-Context Classification
Table 1: Performance—reported in MRR, and when specified in NDCG—of different models on text-to-code with CodeSearchNet using CodeXGLUE. We reported the results presented in CodeT5+, UniXcoder and ModernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024). MoSE surpasses all existing open-source models, achieving an improvement of +3.6% compared to the best-performing open model, CodeT5+. MoSE reduced the performance gap between open-source models and closed-source ones such as OpenAI text-embedding-3-large.
@@
2.2 Self-Distillation: Multi-Layer loss
@@
To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index. The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied. The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set. This Self-Distillation mechanism, shown in Figure 1, allows for flexible model deployment, enabling adaptive layer pruning with minimal impact on performance.
3 Pre-training
We built MoSE on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model. We reduced its original size, obtaining a model of 1B parameters. Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads. MoSE relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1,024 with an intermediate size of 12,288.
To make our model class bidirectional, we removed the causal masking from the self-attention operations within StarCoder-2. Aiming for modularity, we also replaced sliding window attention with full attention. This step was taken to avoid the receptive field phenomenon of the sliding window mechanisms (Zhu et al., 2021). Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference. We have included and summarized additional details about model architecture in appendix D.
3.1 Training Details
We pre-trained MoSE with a batch size of 4M tokens with a maximum context length of 2,048 tokens, for 245,000 training steps, processing approximately 1T tokens from TheStackV2 (Lozhkov et al., 2024) dataset.
We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to $1\mathrm{e}{-6}$, and a weight decay of $1\mathrm{e}{-1}$. We initialized the learning rate at $6.24\mathrm{e}{-4}$ and decreased it using a multi-step learning rate scheduler (Bi et al., 2024) with 4,000 warmup steps. The learning rate was reduced at 120,000, 185,000, 220,000, 230,000, and 240,000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001. We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs, requiring 450,000 GPU working hours on the Leonardo supercomputer (Turisini et al., 2023).
4 Downstream tasks
4.1 Finetuning for Retrieval, text-to-code and code-to-code search